A jet of water with a velocity 30m/s is hitting the wall . Calculate the pressure on the wall .

Grade:10

1 Answers

Arun
6 years ago
Dear student
 
Force of the water exerted on a surface is given byF=ρ*A*V*V
 
Hence pressure = F/A = ρ*V*V = ρ*V²
 
Now pressure = 10³ * 30² = 900 * 10³ = 9*10^5 N/m²
 
 
Regards
Arun (askIITians forum expert)